# Spin-orbit coupling in symmetric and mixed spin-symmetry

### v1

@@ III.2.2 Pair correlation
So far we have shown that the property of the ground state in the three particle system is similar to that in the two particle system regardless of the difference of the spin symmetry. The spatial structure of the ground state discussed in this section rather shows the opposite and reveals that the interplay between SOC and contact interactions brings a distinct feature from the two-particle system.
@@
with $\phi_{i}(x)$ the $i$th eigenstate of the harmonic oscillator and $N=3$ particle number [14]. The pair correlation of the free particle system is given by a two-dimensional Gaussian function, $\rho_{2}(x,y)=|\phi_{0}(x)|^{2}|\phi_{0}(y)|^{2}$. In the limit of strong interactions $g_{\downarrow\uparrow},g\to\infty$, the system corresponds to Tonks-Girardeau gas (TG) gas and behaves as (spinless) fermions [20, 21, 22], and particularly the density profile corresponds to that of the free fermions, $|\Psi_{\text{TG}}(x_{1},x_{2},x_{3})|=|\Psi_{\text{F}}(x_{1},x_{2},x_{3})|$. Therefore, the pair correlation in the limit can be calculated analytically by using the wavefunction of three free fermions trapped in the harmonic potential, given by
@@
III.3 Aligned interaction ($g>0$ and $g_{\downarrow\uparrow}=0$)
Now we turn off anti-aligned interaction $g_{\downarrow\uparrow}=0$ and set non-zero aligned interaction strength $g>0$. Any pseudo-spin configuration of three particle states has the aligned pseudo-spin components and is affected by the aligned interaction. For $\Omega/\omega=0$, the two unaligned pseudo-spin states (the states (iii,iv) mentioned in Sec. III) obtain the lowest energy. For finite $\Omega$, the coherent coupling mixes those states and the two pseudo-spin aligned components $\ket{\downarrow\downarrow\downarrow}$, $\ket{\uparrow\uparrow\uparrow}$. In the limit $\Omega\to\infty$, the contact interaction contribution is not visible, and the ground state approaches that of no interactions.
@@ III.3 Aligned interaction ($g>0$ and $g_{\downarrow\uparrow}=0$)
III.3.1 Pseudo-spin population
@@ III.3.2 Pair correlation
IV Conclusions
We have studied the system of three bosons trapped in a harmonic potential with two internal states in the presence of SOC and investigated the ground state and the energy spectrum by implementing exact diagonalization. Particularly, we focus on how the mixed spin symmetric states contribute to the ground state. We have found that the interplay between the anti-aligned contact interaction and SOC increases the population of the unaligned pseudo-spin components in the ground state. The emergence of the mixed spin symmetric states compensates for the rise of the interaction energy. This is similar to the two-particle system which has different spin symmetry. With the aligned interaction on, the avoided crossing between the ground state and the first excited state is observed only for small interaction, and this causes shape changes in the pseudo-spin populations. However, the emergence of aligned pseudo-spin states does not contribute to the reduction of the interaction energy, and therefore it decays out for strong interaction. Furthermore, we have found that the pair correlation of the ground state shows similarly to the TG gas even for relatively small contact interactions due to the interplay between contact interaction and SOC.
A question arises whether the same behaviour is seen in the more-particle systems. Our results cannot guarantee but implies that this would be the case for the anti-aligned interaction case. In the anti-aligned interaction case, the mixed spin symmetry affects the pseudo-spin populations, and the more-particle systems also have the same spin symmetry. The mixed spin symmetric states may act in the same way.
The mapping between Dicke model and the SOC system with contact interactions was proposed [11, 24, 25], assuming that the pseudo-spin space of the SOC system is only symmetric. However, as shown already, there is significant amount of mixed spin symmetric components even in the ground state for some parameter regime. Therefore, the mapping between these two models are limited.
@@ IV Conclusions
V acknowledgement
Discussions with Thomas Busch, Karol Gietka, Yongping Zhang, Peter Engels, Pere Mujal, and Ian B. Spielman are appreciated. We acknowledge funding from Grant No. PID2020-114626GB-I00 by MCIN/AEI/10.13039/5011 00011033 and “Unit of Excellence María de Maeztu 2020-2023” award to the Institute of Cosmos Sciences, Grant CEX2019-000918-M funded by MCIN/AEI/10.13039/501100011033. We also acknowledge financial support from the Generalitat de Catalunya (Grant 2021SGR01095). A.U acknowledges further support from the Agencia Estatal de Investigación and the Ministerio de Ciencia e Innovación. A.R.-F. acknowledges funding from MIU through Grant No. FPU20/06174.
@@ Appendix A Justification of cut off $M$
In Sec. III, we set the SOC strength $k\xi=4$ and the cut off $M=50$. Here, we show that the cut off $M$ is large enough. As discussed in Sec. II.2, the cut off $M$ is the number of the eigenstates of the harmonic oscillator we take. The SOC couples these eigenstates such that the momentum of up spin or down spin is boosted. Although larger SOC needs larger $M$, the deviation from the results when $k=0$ is negligible if $M$ is large enough.
@@ Appendix B Population of mixed spin symmetry
In this work, we investigate the population of mixed spin symmetry and here explain how to compute it briefly. The mixed spin symmetry basis are described in the first quantization representation in Eq. (II.3). On the other hand, we adopt the second quantization representation for exact diagonalization (see Sec. II.2). Therefore, we transform the second quantization representation of the Hamiltonians (7)(II.2) to the first quantization representation, i.e., label the particles and treat the system as distinguishable particles (see Fig. 9). This enables us to obtain the quantum number and the pseudo-spin state of each particles, i.e., have |$\ket{n_{1},\sigma_{1};n_{2},\sigma_{2};n_{3},\sigma_{3}}$ for $n_{1},n_{2},n_{3}=0,1,\ldots,M-1$ and $\sigma_{1},\sigma_{2},\sigma_{3}=\downarrow,\uparrow$. We transform the pseudo-spin basis using $\downarrow$,$\uparrow$ , ↑ to the classified pseudo-spin basis $S_{j}$, $M_{j}$ by using the relations (II.3)(II.3). For more concrete details, see our code [18].

### v2

@@ IV.2.2 Pair correlation
So far we have shown that properties of the ground state in three particle systems are similar to two particle systems regardless of the difference of the spin symmetry. In this section, we discuss the spatial structure of the ground state and show that the interplay between SOC and contact interactions brings distinct features from the two-particle system.
@@
with $\phi_{i}(x)$ the $i$th eigenstate of the harmonic oscillator and $N=3$ particle number [15]. The pair correlation of non-interacting particles is given by a two-dimensional Gaussian function, $\rho_{2}(x,y)=|\phi_{0}(x)|^{2}|\phi_{0}(y)|^{2}$. In the limit of strong interactions $g_{\downarrow\uparrow},g\to\infty$, the system corresponds to Tonks-Girardeau gas (TG) gas and behaves as (spinless) fermions [21, 22, 23], and particularly the density profile corresponds to that of the free fermions, $|\Psi_{\text{TG}}(x_{1},x_{2},x_{3})|=|\Psi_{\text{F}}(x_{1},x_{2},x_{3})|$. Therefore, the pair correlation in the limit can be calculated analytically by using the wavefunction of three free fermions trapped in the harmonic potential, given by
@@
IV.3 Aligned interaction ($g>0$ and $g_{\downarrow\uparrow}=0$)
Now we turn off anti-aligned interaction $g_{\downarrow\uparrow}=0$ and set non-zero aligned interaction strength $g>0$. Any pseudo-spin configuration of three particle states has the aligned pseudo-spin components and is affected by the aligned interaction. For $\Omega/\omega=0$, the two unaligned pseudo-spin states (the states (iii,iv) mentioned in Sec. IV) obtain the lowest energy. For finite $\Omega$, the coherent coupling mixes those states and the two pseudo-spin aligned components $\ket{\downarrow\downarrow\downarrow}$, $\ket{\uparrow\uparrow\uparrow}$. In the limit $\Omega\to\infty$, the contact interaction contribution is not visible, and the ground state approaches that of no interactions.
@@ IV.3 Aligned interaction ($g>0$ and $g_{\downarrow\uparrow}=0$)
IV.3.1 Pseudo-spin population
@@ IV.3.2 Pair correlation
V Conclusions
We have studied the system of three bosons trapped in a harmonic potential with two internal states in the presence of SOC and investigated the ground state and the energy spectrum by implementing exact diagonalization. Particularly, we have focused on how the mixed spin symmetric states contribute to the ground state and have found similarity and difference. We have found that the interplay between the anti-aligned contact interaction and SOC increases the population of the unaligned pseudo-spin components in the ground state. This is due to the emergence of the mixed spin symmetric states, which can be partially anti-symmetric. This anti-symmetry compensates for the rise of the interaction energy. Moreover, with the aligned interactions, the avoided crossing between the ground state and the first excited state is observed only for small interaction, and this causes shape changes in the pseudo-spin populations. However, the emergence of aligned pseudo-spin states does not contribute to the reduction of the interaction energy, and therefore it decays out for strong interaction. This is similar to the two-particle system. In addition, we have found that the pair correlation of the ground state has a structure similar to the TG gas even for relatively small contact interactions. This strong interaction effect is not seen in two-particle systems and is interesting because there are only aligned or anti-aligned interactions.
A natural question is whether the same behaviour is seen in the more-particle systems. Our results cannot guarantee but imply that this would be the case for anti-aligned interactions. In these cases, the mixed spin symmetry affects the pseudo-spin populations, and more-particle systems also have the same spin symmetry. The mixed spin symmetric states may act in the same way. Furthermore, we have studied only contact interactions. The fact that the anti-symmetric pseudo-spin states do not feel the interaction depends on the kind of the interaction, and the results we have shown may change with different interaction types.
The mapping between Dicke model and SOC systems with contact interactions was proposed [11, 25, 26], assuming that the pseudo-spin space of the SOC system is only symmetric. However, as shown already, there is significant amount of mixed spin symmetric components even in the ground state for some parameter regime. Therefore, the mapping between these two models is limited.
@@ V Conclusions
VI acknowledgement
Discussions with Thomas Busch, Karol Gietka, Yongping Zhang, Peter Engels, Pere Mujal, and Ian B. Spielman are gratefully appreciated. We acknowledge funding from Grant No. PID2020-114626GB-I00 by MCIN/AEI/10.13039/5011 00011033 and “Unit of Excellence María de Maeztu 2020-2023” award to the Institute of Cosmos Sciences, Grant CEX2019-000918-M funded by MCIN/AEI/10.13039/501100011033. We also acknowledge financial support from the Generalitat de Catalunya (Grant 2021SGR01095). A.U acknowledges further support from the Agencia Estatal de Investigación and the Ministerio de Ciencia e Innovación. A.R.-F. acknowledges funding from MIU through Grant No. FPU20/06174.
@@ Appendix A Justification of cut off $M$
In Sec. IV, we set the SOC strength $k\xi=4$ and the cut off $M=50$. Here, we show that the cut off $M$ is large enough. As discussed in Sec. III.2, the cut off $M$ is the number of the eigenstates of the harmonic oscillator we take. The SOC couples these eigenstates such that the momentum of up spin or down spin is boosted. Although larger SOC needs larger $M$, the deviation from the results when $k=0$ is negligible if $M$ is large enough.
@@ Appendix B Population of mixed spin symmetry
In this work, we investigate the population of mixed spin symmetry and here explain how to compute it briefly. The mixed spin symmetry basis are described in the first quantization representation in Eq. (II). On the other hand, we adopt the second quantization representation for exact diagonalization (see Sec. III.2). Therefore, we transform the second quantization representation of the Hamiltonians (7)(III.2) to the first quantization representation, i.e., label the particles and treat the system as distinguishable particles (see Fig. 9). This enables us to obtain the quantum number and the pseudo-spin state of each particles, i.e., have |$\ket{n_{1},\sigma_{1};n_{2},\sigma_{2};n_{3},\sigma_{3}}$ for $n_{1},n_{2},n_{3}=0,1,\ldots,M-1$ and $\sigma_{1},\sigma_{2},\sigma_{3}=\downarrow,\uparrow$. We transform the pseudo-spin basis using $\downarrow$,$\uparrow$ , ↑ to the classified pseudo-spin basis $S_{j}$, $M_{j}$ by using the relations (II)(II). For more concrete details, see our code [19].